Q:-
Before 5 years, A father is 3 time older than his son.
after 10 year he becomes twice.
what is the present age of the father and son?.

Answer by lwsshak3(11628)

Before 5 years, A father is 3 time older than his son.
after 10 year he becomes twice.
what is the present age of the father and son?.
***
Before 5 years:
x=son's age
3x=father's age
..
At present:
x+5=age of son.
3x+5=age of father.
..
After 10 years:
x+15=son's age
3x+15=father's age
3x+15=2(x+15)
3x+15=2x+30
x=15
..
x+5=20
3x+5=50
..
present age of son=20
present age of father=50